OpenStudy (anonymous):

In a dog show, 20% of the dogs are eliminated each day. If the competition starts with 2,000 dogs, to the nearest integer, how many dogs will be present at the start of the 5th day?

OpenStudy (anonymous):

4 is for the days.. 20/100 is for 20%

OpenStudy (anonymous):

are you sure its not 819?

OpenStudy (anonymous):

i think you are supposed to do this \[2000\times (.8)^4\]

OpenStudy (anonymous):

day one 2000 day two \(2000 \times .8\) day three \(2000\times .8^2\) day four \(2000\times .8^3\) day five \(2000\times .8^4\) reasoning that reducing a number by 20% is the same as multiplying by .8

OpenStudy (anonymous):

i get 819 using that method, yes

OpenStudy (anonymous):

aha.. then I'm sorry @meowmeow345 :)